OpenStudy (anonymous):

what is 90% of 250? which equation could be used to answer this question? A. x=0.9 x 250 B. 90=250x C. x=0.09 x 250 D. 250 = x . 90

OpenStudy (anonymous):

do you have a guess on which?

OpenStudy (anonymous):

x = (90/100)*250 x = (90*250)/100 x = 9*25 Multiply and get the solution..

OpenStudy (anonymous):

is it 225

OpenStudy (anonymous):

Sure is.

OpenStudy (anonymous):

You can check your answer by doing 225/250 = .9 which is 90%

OpenStudy (anonymous):

so is it a b c or d

OpenStudy (anonymous):

which one gives you 225?

OpenStudy (anonymous):

I think this is about finding x right? so i think choose a = 0.9*250

OpenStudy (anonymous):

225 a